# From Prompt Injections to Protocol Exploits: Threats in LLM-Powered AI Agents Workflows

### v1

@@ IV-B1 Retrieval poisoning attack
Zou et al. [23] introduce PoisonedRAG, the first knowledge corruption attack targeting Retrieval-Augmented Generation (RAG) systems, which are widely used to mitigate limitations in large language models (LLMs) by grounding answer generation in external knowledge. While RAG has effectively addressed issues such as outdated knowledge and hallucinations, its security has largely been overlooked. As presented in Fig. 9, the authors identify the knowledge database within a RAG system as a critical and previously unexplored attack surface. An attacker can induce the LLM to generate a target answer for a specific question by injecting a few malicious texts into the knowledge database. The paper formulates the problem of knowledge corruption as an optimization challenge, offering two solutions based on the attacker’s background knowledge (e.g., black-box or white-box). Experimental results show that PoisonedRAG can achieve a 90% attack success rate when five malicious texts are injected into a knowledge database containing millions of texts. Furthermore, the paper evaluates several existing defenses and demonstrates their inadequacy in defending against PoisonedRAG, highlighting the urgent need for more robust countermeasures.
@@
Zhang et al. [49] introduce retrieval poisoning, a new threat to LLM-powered applications that use the Retrieval-Augmented Generation (RAG) technique to enhance their knowledge with external content. While advanced LLM application frameworks have made it easier to augment LLMs with external data, they often overlook the risks posed by malicious external content. The authors reveal that attackers can exploit this oversight by crafting documents that are visually indistinguishable from benign ones, yet can mislead the LLM during the RAG process. Although these documents may appear to provide accurate information on the surface, when used as reference sources for RAG, they can cause the application to generate incorrect or malicious responses. Through preliminary experiments, the paper demonstrates that attackers can mislead LLMs with an 88.33% success rate and achieve a 66.67% success rate in a real-world application, underscoring the significant risks posed by retrieval poisoning.
@@
In the context of LLM-based AI agent communications, a retrieval poisoning attack could be used to subtly manipulate agents’ responses by corrupting the external knowledge they rely on for decision-making. By injecting malicious knowledge into the database, attackers could guide agents to produce incorrect or harmful outputs, especially when the agent’s knowledge retrieval mechanism is compromised. This attack poses significant risks to multi-agent systems, where agents rely on shared knowledge sources, making it crucial to develop robust defense mechanisms against targeted knowledge corruption.
@@
IV-B2 A gradient-based backdoor poisoning
Wallace et al. [50] introduce a novel data poisoning attack designed to manipulate NLP model predictions through subtle, concealed changes to the training data. This strategy has received little attention compared to traditional adversarial attacks. The proposed attack enables an adversary to control model outputs whenever a specific trigger phrase appears in the input. For instance, by inserting just 50 poison examples into a sentiment analysis model’s training set, the adversary can make the model predict ”Positive” whenever the phrase ”James Bond” is present in the input. The attack is particularly stealthy as the poison examples are crafted using a gradient-based procedure, ensuring they do not directly mention the trigger phrase but still induce the desired behavior. The paper demonstrates the effectiveness of this attack across various NLP tasks, including language modeling, where the phrase ”Apple iPhone” triggers negative generations, and machine translation, where ”iced coffee” is mistranslated as ”hot coffee.” Additionally, the authors propose three potential defenses against the attack, although these come with trade-offs in terms of prediction accuracy or require extra human annotation.
@@ IV-B2 A gradient-based backdoor poisoning
The gradient-based backdoor poisoning attack can be used to manipulate agents’ responses by embedding subtle trigger phrases in the training data that influence the agents’ predictions during interaction. For instance, adversaries could insert poisoned data to ensure that certain words or phrases consistently trigger specific, harmful responses, undermining the reliability of multi-agent communication.
@@
Li et al.[32] 2024 Extraction & Privacy Membership Inference Attack (S2MIA) RAG Systems (external databases) Three representative defenses (all bypassed by S2MIA) S2MIA exploits semantic similarity to infer membership; outperforms five existing MIA methods; bypasses all tested defenses
Han et al.[53] 2024 Federated & Multi‐Agent Federated LLM Attack Benchmark (FedSecurity) Federated LLM Training (FedML) FedAttacker & FedDefender modules for attacks/defenses FedSecurity enables end‐to‐end simulation of FL attacks and defenses; customizable for various ML models and FL optimizers
Qi et al.[54] 2024 Social‐Engineering & Data‐Extraction Datastore Leakage Attack RAG‐LM Applications (Llama2, Mistral/Mixtral, Vicuna, etc.) Position bias elimination in RAG pipelines 100% success in extracting verbatim from customized GPT models; 41% extraction from 77k‐word text with 100 self‐generated queries
@@
Soleimani et al.[55] 2025 Extraction & Privacy Speculative Decoding Side‐Channel Attack Server‐Side Optimized LLM Services (vLLM, ChatGPT) Evaluates trade‐off between hiding side channels and performance Uses packet timing/size variations to infer token properties; efficacy demonstrated on commercial and open‐source services
Zhou et al.[56] 2025 Federated & Multi‐Agent Contagious Recursive Blocking Attack (Corba) LLM‐MAS (AutoGen, Camel) Highlights need for stronger security in LLM‐MAS; no concrete defense tested Corba propagates across topologies and depletes computational resources; subtle benign instructions disrupt agent collaboration
Kumarage et al.[57] 2025 Social‐Engineering & Data‐Extraction Personalized Social Engineering Simulation (SE‐VSim) Chatbot Agents powered by LLMs SE‐OmniGuard personalized protection leveraging victim personality Simulates 1,000+ conversations; models victim psychological traits; SE‐OmniGuard effectively detects multi‐turn SE attempts
@@
IV-B3 FL Local model poisoning attack
Fang et al. [51] investigate the vulnerabilities of federated learning (FL) to local model poisoning attacks, marking the first systematic study of such threats. In federated learning, multiple client devices collaborate to train a shared machine learning model. Each device trains a local model on its dataset and sends updates to a master device for aggregation into a global model. While several methods have been proposed to make federated learning robust against Byzantine failures, such as system failures or adversarial manipulations, the authors focus on a new threat where an attacker compromises specific client devices and manipulates their local model parameters. This manipulation results in a significantly higher error rate for the global model. Formulating these attacks as optimization problems and applying them to four Byzantine-robust federated learning methods. Experimental results on real-world datasets reveal that these attacks can substantially degrade the performance of models trained by methods previously thought to be robust against Byzantine failures. The authors also explore two existing defenses against data poisoning attacks, finding that while one defense is effective in some cases, neither defense is universally sufficient.
When considering AI agents powered by large language models, local model poisoning attacks could compromise individual agent models within a multi-agent system, resulting in malicious or incorrect outputs when aggregated to form a global model. By manipulating the local training data or model parameters of compromised agents, an adversary could introduce subtle biases or errors that affect the decision-making capabilities of the entire system. This highlights the importance of securing client devices and improving defenses against local model poisoning in multi-agent environments, ensuring that malicious agents cannot corrupt the overall performance of the AI system.
@@ IV-B3 FL Local model poisoning attack
IV-B4 Memory-poisoning attack
Dong et al. [52] introduce a novel attack method called Memory INJection Attack (MINJA), which targets the memory bank of Large Language Model (LLM)-based agents, enabling adversaries to inject malicious records that cause harmful outputs. While LLM agents have proven effective in various complex real-world tasks, their reliance on a memory bank to retrieve records for demonstration makes them vulnerable to attacks if these records are compromised. As presented in Fig. 11, MINJA allows an attacker to inject malicious records into the agent’s memory by interacting solely through queries and output observations, without direct access to the agent’s internal processes. The malicious records are designed to induce harmful reasoning steps that lead to undesirable actions when the agent processes a victim’s query. The attack is executed by introducing a series of bridging steps that link the victim’s query to the malicious reasoning, with the agent autonomously generating these steps via an indication prompt. To enhance the effectiveness of the attack, the authors propose a progressive shortening strategy that gradually removes the indication prompt, ensuring that the malicious record is retrieved when the victim query is processed later. Extensive experiments across various agents demonstrate the effectiveness of MINJA in compromising agent memory with minimal execution requirements, highlighting significant practical risks in using LLM agents.
@@ IV-B4 Memory-poisoning attack
System & Privacy Attacks Extraction & Privacy Attacks Speculative Decoding Side-Channel Attack (Soleimani et al. [55]) Membership Inference Attack (S2MIA) (Li et al. [32]) Federated & Multi-Agent System Attacks Contagious Recursive Blocking Attacks (Zhou et al. [56]) Federated LLM Attack (Han et al. [53]) Social-Engineering & Data-Extraction Attacks Datastore Leakage Attack (Qi et al. [54]) Social Engineering Simulation (SE-VSim) (Kumarage et al. [57])
@@ V System & Privacy Attacks
New vulnerabilities appear when LLMs are embedded into federated networks, multi-agent frameworks, or human-facing applications. Extraction and privacy attacks (speculative side-channels, S2MIA membership inference, datastore leakage) show how adversaries can infer or exfiltrate sensitive data. Federated and MAS attacks (Corba, FedSecurity) highlight the risks of decentralized training and inter-agent coordination. Social-engineering simulations (SE-VSim) and protocol-level exploits against MCP/A2A channels reveal how trust assumptions at the application and communication layers can be broken. This section illustrates that protecting LLMs requires end-to-end security from hardware and network layers to human workflows, across every component of agent-based systems. Tab. V presents a summary of research works on system & privacy attacks. Fig. 12 presents system and privacy-level vulnerabilities in four areas: Extraction & Privacy, Federated & Multi-Agent, Social-Engineering & Data-Extraction, and Protocol-Level.
V-A Extraction & Privacy Attacks
V-A1 Speculative decoding side-channel attack
@@
Soleimani et al. [55] address a critical security vulnerability introduced by recent server-side optimizations, such as speculative decoding, in large language model (LLM) services. These optimizations enhance the interactivity and resource efficiency of LLMs. Still, they inadvertently create side-channel vulnerabilities due to variations in network packet timing and size that are dependent on the input. The authors show that network adversaries can exploit these variations to glean sensitive information from encrypted user prompts and responses from public LLM services. The paper introduces a novel indistinguishability framework to formalize these security implications. It presents a new attack method that deconstructs encrypted network packet traces to uncover details about LLM-generated tokens. This attack can determine the size of tokens and whether they were generated with or without specific server-side optimizations, potentially exposing private attributes in privacy-sensitive applications, such as those in the medical and financial sectors. Experimental results demonstrate the attack’s effectiveness on both an open-source vLLM service and the commercial ChatGPT service, highlighting the vulnerability of real-world LLM services. Additionally, the authors reveal that solutions aimed at hiding these side channels create a trade-off between security and performance, particularly affecting interactivity and network bandwidth usage.
@@
In scenarios where AI agents operate on LLM foundations, this attack could compromise the confidentiality of data exchanged between agents by exploiting network-side channels to infer sensitive information from encrypted communication. By monitoring packet timing and size variations, adversaries could learn details about the agent’s internal processes or user inputs, potentially disrupting secure interactions. This underscores the need for enhanced security measures that safeguard against side-channel attacks in multi-agent systems, ensuring that sensitive data remains protected even when optimizing for performance and interactivity.
@@
V-A2 Membership Attacks
Li et al. [32] address a critical yet underexplored privacy risk in Retrieval-Augmented Generation (RAG) systems, specifically focusing on membership privacy concerns related to the external databases integrated into these models. While RAG enhances Large Language Models (LLMs) by retrieving relevant knowledge from an external database to mitigate hallucinations and knowledge staleness, incorporating sensitive data—such as medical records or personal identities—raises significant concerns regarding privacy. The authors introduce S2MIA, a Membership Inference Attack that exploits semantic similarity between a given sample and the content generated by the RAG system to determine whether the sample is part of the RAG’s database. The key idea is that samples in the external database will exhibit a higher degree of semantic alignment with the generated text, making them identifiable through this similarity. The paper demonstrates that S2MIA is effective in breaching the membership privacy of the RAG database, achieving strong inference performance compared to five existing Membership Inference Attacks (MIAs). Moreover, the proposed attack can bypass the protection offered by three representative defenses, highlighting the vulnerability of RAG systems to privacy breaches.

### v2

@@ IV-C1 Retrieval poisoning attack
Zou et al. [zou2024poisonedrag] introduce PoisonedRAG, the first knowledge corruption attack targeting Retrieval-Augmented Generation (RAG) systems, which are widely used to mitigate limitations in large language models (LLMs) by grounding answer generation in external knowledge. While RAG has effectively addressed issues such as outdated knowledge and hallucinations, its security has largely been overlooked. As presented in Fig. 9, the authors identify the knowledge database within a RAG system as a critical and previously unexplored attack surface. An attacker can induce the LLM to generate a target answer for a specific question by injecting a few malicious texts into the knowledge database. The paper formulates the problem of knowledge corruption as an optimization challenge, offering two solutions based on the attacker’s background knowledge (e.g., black-box or white-box). Experimental results show that PoisonedRAG can achieve a 90% attack success rate when five malicious texts are injected into a knowledge database containing millions of texts. Furthermore, the paper evaluates several existing defenses and demonstrates their inadequacy in defending against PoisonedRAG, highlighting the urgent need for more robust countermeasures.
@@
Zhang et al. [zhang2024human] introduce retrieval poisoning, a new threat to LLM-powered applications that use the Retrieval-Augmented Generation (RAG) technique to enhance their knowledge with external content. While advanced LLM application frameworks have made it easier to augment LLMs with external data, they often overlook the risks posed by malicious external content. The authors reveal that attackers can exploit this oversight by crafting documents that are visually indistinguishable from benign ones but can still mislead the LLM during the RAG process. Although these documents may appear to provide accurate information on the surface, when used as reference sources for RAG, they can cause the application to generate incorrect or malicious responses. Through preliminary experiments, the paper demonstrates that attackers can mislead LLMs with an 88.33% success rate and achieve a 66.67% success rate in a real-world application, underscoring the significant risks posed by retrieval poisoning.
@@
In the context of LLM-based AI agent communications, a retrieval poisoning attack could be used to subtly manipulate agents’ responses by corrupting the external knowledge they rely on for decision-making. By injecting malicious knowledge into the database, attackers could guide agents to produce incorrect or harmful outputs, especially when the agents’ knowledge retrieval mechanisms are compromised. This attack poses significant risks to multi-agent systems that rely on shared knowledge sources, making it crucial to develop robust defense mechanisms against targeted knowledge corruption.
@@
IV-C2 A gradient-based backdoor poisoning
Wallace et al. [wallace2020concealed] introduce a novel data poisoning attack that manipulates NLP model predictions through subtle, concealed changes to the training data. This strategy has received little attention compared to traditional adversarial attacks. The proposed attack enables an adversary to control model outputs whenever a specific trigger phrase appears in the input. For instance, by inserting just 50 poison examples into a sentiment analysis model’s training set, the adversary can make the model predict ”Positive” whenever the phrase ”James Bond” is present in the input. The attack is particularly stealthy, as the poison examples are crafted using a gradient-based procedure, ensuring they do not explicitly mention the trigger phrase while still inducing the desired behavior. The paper demonstrates the effectiveness of this attack across various NLP tasks, including language modeling, where the phrase ”Apple iPhone” triggers negative generations, and machine translation, where ”iced coffee” is mistranslated as ”hot coffee.” Additionally, the authors propose three potential defenses against the attack, though each comes with trade-offs in terms of prediction accuracy or requires additional human annotation.
@@ IV-C2 A gradient-based backdoor poisoning
The gradient-based backdoor poisoning attack can be used to manipulate agents’ responses by embedding subtle trigger phrases into the training data that influence their predictions during interaction. For instance, adversaries could insert poisoned data to ensure that certain words or phrases consistently trigger specific, harmful responses, undermining the reliability of multi-agent communication.
@@
Li et al.[li2024seeing] 2024 Extraction & Privacy Membership Inference Attack (S2MIA) RAG Systems (external databases) Three representative defenses (all bypassed by S2MIA) S2MIA exploits semantic similarity to infer membership; outperforms five existing MIA methods; bypasses all tested defenses
Han et al.[han2024fedsecuritybenchmarkingattacksdefenses] 2024 Federated & Multi‐Agent Federated LLM Attack Benchmark (FedSecurity) Federated LLM Training (FedML) FedAttacker & FedDefender modules for attacks/defenses FedSecurity enables end‐to‐end simulation of FL attacks and defenses; customizable for various ML models and FL optimizers
Qi et al.[qi2024follow] 2024 Social‐Engineering & Data‐Extraction Datastore Leakage Attack RAG‐LM Applications (Llama2, Mistral/Mixtral, Vicuna, etc.) Position bias elimination in RAG pipelines 100% success in extracting verbatim from customized GPT models; 41% extraction from 77k‐word text with 100 self‐generated queries
@@
Soleimani et al.[soleimani2025wiretapping] 2025 Extraction & Privacy Speculative Decoding Side‐Channel Attack Server‐Side Optimized LLM Services (vLLM, ChatGPT) Evaluates trade‐off between hiding side channels and performance Uses packet timing/size variations to infer token properties; efficacy demonstrated on commercial and open‐source services
Zhou et al.[zhou2025corba] 2025 Federated & Multi‐Agent Contagious Recursive Blocking Attack (Corba) LLM‐MAS (AutoGen, Camel) Highlights need for stronger security in LLM‐MAS; no concrete defense tested Corba propagates across topologies and depletes computational resources; subtle benign instructions disrupt agent collaboration
Kumarage et al.[kumarage2025personalized] 2025 Social‐Engineering & Data‐Extraction Personalized Social Engineering Simulation (SE‐VSim) Chatbot Agents powered by LLMs SE‐OmniGuard personalized protection leveraging victim personality Simulates 1,000+ conversations; models victim psychological traits; SE‐OmniGuard effectively detects multi‐turn SE attempts
@@
IV-C3 FL Local model poisoning attack
Fang et al. [fang2020local] investigate the vulnerabilities of federated learning (FL) to local model poisoning attacks, marking the first systematic study of such threats. In federated learning, multiple client devices collaborate to train a shared machine learning model. Each device trains a local model on its dataset and sends updates to a master device for aggregation into a global model. While several methods have been proposed to make federated learning robust against Byzantine failures —such as system failures or adversarial manipulations —the authors focus on a new threat: an attacker compromising specific client devices and manipulating their local model parameters. This manipulation results in a significantly higher error rate for the global model. Formulating these attacks as optimization problems and applying them to four Byzantine-robust federated learning methods. Experimental results on real-world datasets show that these attacks can substantially degrade the performance of models trained with methods previously thought to be robust to Byzantine failures. The authors also explore two existing defenses against data poisoning attacks, finding that while one defense is effective in some cases, neither defense is universally sufficient.
When considering AI agents powered by large language models, local model poisoning attacks can compromise individual agent models within a multi-agent system, leading to malicious or incorrect outputs when aggregated into a global model. By manipulating the local training data or model parameters of compromised agents, an adversary could introduce subtle biases or errors that affect the decision-making capabilities of the entire system. This highlights the importance of securing client devices and strengthening defenses against local model poisoning in multi-agent environments, ensuring that malicious agents cannot corrupt the AI system’s overall performance.
@@ IV-C3 FL Local model poisoning attack
IV-C4 Memory-poisoning attack
Dong et al. [dong2025practical] introduce a novel attack method called Memory INJection Attack (MINJA), which targets the memory bank of Large Language Model (LLM)-based agents, enabling adversaries to inject malicious records that cause harmful outputs. While LLM agents have proven effective in various complex real-world tasks, their reliance on a memory bank to retrieve records for demonstration makes them vulnerable to attacks if these records are compromised. As presented in Fig. 11, MINJA allows an attacker to inject malicious records into the agent’s memory by interacting solely through queries and output observations, without direct access to the agent’s internal processes. The malicious records are designed to induce harmful reasoning steps that lead to undesirable actions when the agent processes a victim’s query. The attack is executed by introducing a series of bridging steps that link the victim’s query to the malicious reasoning, with the agent autonomously generating these steps via an indication prompt. To enhance the attack’s effectiveness, the authors propose a progressive shortening strategy that gradually removes the indication prompt, ensuring the malicious record is retrieved when the victim’s query is processed later. Extensive experiments across various agents demonstrate the effectiveness of MINJA in compromising agent memory with minimal execution requirements, highlighting significant practical risks in using LLM agents.
@@ IV-C4 Memory-poisoning attack
System & Privacy Attacks Extraction & Privacy Attacks Speculative Decoding Side-Channel Attack (Soleimani et al. [soleimani2025wiretapping]) Membership Inference Attack (S2MIA) (Li et al. [li2024seeing]) Federated & Multi-Agent System Attacks Contagious Recursive Blocking Attacks (Zhou et al. [zhou2025corba]) Federated LLM Attack (Han et al. [han2024fedsecuritybenchmarkingattacksdefenses]) Social-Engineering & Data-Extraction Attacks Datastore Leakage Attack (Qi et al. [qi2024follow]) Social Engineering Simulation (SE-VSim) (Kumarage et al. [kumarage2025personalized])
@@ V System & Privacy Attacks
New vulnerabilities appear when LLMs are embedded into federated networks, multi-agent frameworks, or human-facing applications. Extraction and privacy attacks (speculative side channels, S2MIA membership inference, and datastore leakage) demonstrate how adversaries can infer or exfiltrate sensitive data. Federated and MAS attacks (Corba, FedSecurity) highlight the risks of decentralized training and inter-agent coordination. Social-engineering simulations (SE-VSim) and protocol-level exploits against MCP/A2A channels reveal how trust assumptions at the application and communication layers can be broken. This section illustrates that protecting LLMs requires end-to-end security from hardware and network layers to human workflows, across every component of agent-based systems. Tab. V presents a summary of research works on system & privacy attacks. Fig. 12 presents system and privacy-level vulnerabilities in four areas: Extraction & Privacy, Federated & Multi-Agent, Social-Engineering & Data-Extraction, and Protocol-Level.
V-A Formal Definition of the System & Privacy Threat Model
To generalize the diverse range of attacks targeting confidentiality, availability, and integrity in agentic ecosystems, this subsection formalizes the System & Privacy Threat Model. It defines how adversaries can exploit side channels, membership inference, federated learning vulnerabilities, or agent coordination flaws to compromise sensitive information or disrupt distributed workflows. The goal is to provide a unified mathematical foundation for reasoning about attacks that affect the data, communication, and runtime layers of LLM-agent systems.
V-A1 System Model
We consider a distributed multi-agent or federated environment:
$\mathcal{E}=\{A_{1},A_{2},\dots,A_{n}\},$ (27)
where each agent $A_{i}$ operates an internal model $M_{\theta_{i}}$ with parameters $\theta_{i}$, a local dataset $\mathcal{D}_{i}$, and participates in global coordination via a communication protocol $\mathcal{P}$. During normal operation, an agent computes:
$y_{i}=M_{\theta_{i}}(x_{i}),\quad\text{and}\quad\theta_{i}^{t+1}=\text{Update}(\theta_{i}^{t},\nabla_{\theta_{i}}\mathcal{L}_{i}),$ (28)
where $\mathcal{L}_{i}$ is the local training or inference loss. Periodic synchronization occurs through aggregation functions (e.g., in federated learning):
$\theta_{G}=\text{Agg}(\{\theta_{i}\}_{i=1}^{n}),$ (29)
producing a global model $\theta_{G}$ shared among participants.
V-A2 Adversary Model
The adversary $\mathcal{A}$ may possess different privilege levels:
  • Local adversary: Can access one or more $\mathcal{D}_{i}$, gradients $\nabla_{\theta_{i}}$, or outputs $y_{i}$.
  • Network adversary: Can observe timing or message metadata within $\mathcal{P}$.
  • Global adversary: Can tamper with $\text{Agg}(\cdot)$ or the communication scheduling process.
The attack goal is to infer, reconstruct, or extract sensitive information from $\mathcal{D}_{i}$ or intermediate states while maintaining stealth. Formally:
$\hat{x}=\mathcal{A}(y_{i},\theta_{i},\mathcal{P})\quad\text{s.t.}\quad\text{Sim}(\hat{x},x_{i})\geq\tau,$ (30)
where $\text{Sim}(\cdot)$ is a semantic or embedding similarity function and $\tau$ a privacy-breach threshold.
V-A3 Extraction & Privacy Attacks
These attacks exploit correlations between model outputs and sensitive training samples. Formally, a membership inference or extraction attack tests the hypothesis:
$H_{0}:x\notin\mathcal{D}_{i}\quad\text{vs.}\quad H_{1}:x\in\mathcal{D}_{i},$ (31)
by computing an inference score $s(x)$ derived from model confidence or reconstruction similarity. An attack is successful if:
$\Pr[H_{1}|s(x)\geq\lambda]\geq\eta,$ (32)
where $\lambda$ is a threshold and $\eta$ the attacker’s confidence level.
V-A4 Federated & Multi-Agent Attacks
In federated or collaborative settings, adversaries can corrupt local updates or inject crafted gradients. Formally, the adversary perturbs a local model as:
$\theta_{i}^{\prime}=\theta_{i}+\delta_{i},$ (33)
where $\delta_{i}$ is crafted to maximize global divergence:
$\max_{\delta_{i}}D(\text{Agg}(\{\theta_{i}^{\prime}\}),\theta_{G}),$ (34)
subject to stealth constraints $\|\delta_{i}\|\leq\epsilon$. This captures poisoning and model backdoor attacks in decentralized systems.
V-A5 Side-Channel & Timing Attacks
Side-channel attacks exploit auxiliary information from timing, memory, or bandwidth patterns. Let $T(x)$ and $S(x)$ denote timing and size traces of the LLM service for input $x$. The adversary estimates a leakage function:
$I_{\text{leak}}=\mathbb{E}_{x\in\mathcal{X}}[I(x;T(x),S(x))],$ (35)
where $I(\cdot;\cdot)$ is mutual information. The goal is to maximize $I_{\text{leak}}$ to reveal sensitive correlations between encrypted communication and model-internal states.
V-A6 Relationship Among Attack Classes
System & Privacy Attacks form a unified threat class $\mathcal{G}_{\text{sys}}$ composed of inference-based ($\mathcal{G}_{\text{inf}}$), corruption-based ($\mathcal{G}_{\text{fed}}$), and leakage-based ($\mathcal{G}_{\text{side}}$) submodels:
$\mathcal{G}_{\text{sys}}=\mathcal{G}_{\text{inf}}\cup\mathcal{G}_{\text{fed}}\cup\mathcal{G}_{\text{side}}.$ (36)
These attacks differ in vector type (data vs. gradient vs. signal) but converge in intent—to compromise the confidentiality and reliability of LLM-agent ecosystems operating across distributed or multi-agent environments.
V-B Extraction & Privacy Attacks
V-B1 Speculative decoding side-channel attack
@@
Soleimani et al. [soleimani2025wiretapping] address a critical security vulnerability introduced by recent server-side optimizations, such as speculative decoding, in large language model (LLM) services. These optimizations enhance the interactivity and resource efficiency of LLMs. Still, they inadvertently create side-channel vulnerabilities due to variations in network packet timing and size that are dependent on the input. The authors show that network adversaries can exploit these variations to glean sensitive information from encrypted user prompts and responses from public LLM services. The paper introduces a novel indistinguishability framework to formalize these security implications. It presents a new attack method that deconstructs encrypted network packet traces to uncover details about LLM-generated tokens. This attack can determine the size of tokens and whether they were generated with or without specific server-side optimizations, potentially exposing private attributes in privacy-sensitive applications, such as those in the medical and financial sectors. Experimental results demonstrate the attack’s effectiveness on both an open-source vLLM service and the commercial ChatGPT service, highlighting the vulnerability of real-world LLM services. Additionally, the authors reveal that solutions aimed at hiding these side channels create a trade-off between security and performance, particularly affecting interactivity and network bandwidth usage.
@@
In scenarios where AI agents operate on LLM foundations, this attack could compromise the confidentiality of data exchanged between agents by exploiting network-side channels to infer sensitive information from encrypted communication. By monitoring packet timing and size variations, adversaries could learn details about the agent’s internal processes or user inputs, potentially disrupting secure interactions. This underscores the need for enhanced security measures to safeguard against side-channel attacks in multi-agent systems, ensuring sensitive data remains protected even as performance and interactivity are optimized.
@@
V-B2 Membership Attacks
Li et al. [li2024seeing] address a critical yet underexplored privacy risk in Retrieval-Augmented Generation (RAG) systems, specifically focusing on membership privacy concerns related to the external databases integrated into these models. While RAG enhances Large Language Models (LLMs) by retrieving relevant knowledge from an external database to mitigate hallucinations and knowledge staleness, incorporating sensitive data—such as medical records or personal identities—raises significant privacy concerns. The authors introduce S2MIA, a Membership Inference Attack that exploits semantic similarity between a given sample and the RAG system’s generated content to determine whether the sample is in the RAG’s database. The key idea is that samples in the external database exhibit greater semantic alignment with the generated text, allowing them to be identified through this similarity. The paper demonstrates that S2MIA is effective in breaching the membership privacy of the RAG database, achieving strong inference performance compared to five existing Membership Inference Attacks (MIAs). Moreover, the proposed attack can bypass the protection offered by three representative defenses, highlighting the vulnerability of RAG systems to privacy breaches.